### v1

# Optimum Spectrum Extension for PAPR Reduction of DFT-s-OFDM

Abstract
Uplink coverage in cellular networks is constrained by the maximum UE transmit power, making peak-to-average power ratio (PAPR) reduction essential. While DFT-s-OFDM with frequency-domain spectral shaping (FDSS) achieves significantly lower PAPR than OFDM, especially with $\pi/2$-BPSK, the PAPR remains too high for higher-rate transmission. Spectrum extension (SE) combined with FDSS (FDSS-SE) can further reduce the PAPR for higher-order QAM. This paper considers FDSS-SE with parametrized FDSS windows spanning a range of possible power ripples, as well as arbitrary circular shifts of the subcarrier coefficients. We optimize both the frequency shift and the SE size, and show that there exists an optimal SE size for reducing the PAPR and another one for increasing the rate. Analysis and simulations reveal that both optima largely depend on the window attenuation but are nearly invariant in proportion to the bandwidth. While the PAPR-optimal SE size is nearly invariant to the constellation order of regular QAM, the rate-optimal SE size depends also on the SNR. These insights provide practical guidelines for beyond-5G uplink coverage enhancement, highlighting that SE size should be individually configured according to the user’s FDSS window and link quality.
I Introduction
Uplink transmission is typically the coverage bottleneck of modern wireless cellular systems due to the limited transmission power of user equipment (UE). As a result, uplink coverage enhancement remains an on-going topic in 3GPP standardization, as it has been for 5G [1, 2] and expected to continue into 6G [3].
In coverage-limited scenarios, transmitting at the highest possible average power is desirable to maximize the link quality. However, the dynamic range of the transmitted signals, commonly measured by its peak-to-average power ratio (PAPR), often forces the UE to back off from its maximum power to avoid distortion from peaks penetrating the non-linear region of the power amplifier. This back-off prevents uncontrolled distortion but directly reduces the signal-to-noise ratio (SNR).
Orthogonal frequency division multiplexing (OFDM) is the dominant modulating waveform of wireless communications and forms the basis of 3GPP standards for both uplink and downlink. Discrete Fourier transform-spread OFDM (DFT-s-OFDM) is an uplink option in 5G NR that reduces signal fluctuations compared to conventional OFDM. In DFT-s-OFDM, data symbols undergo DFT precoding before OFDM modulation, producing a single-carrier-like waveform since each modulation symbol’s energy is distributed across all allocated subcarriers. Reciprocally, in the time domain, a large fraction of a symbol’s energy is concentrated in a short interval, leading to a pulse-multiplexing of the data. While DFT-s-OFDM improves the PAPR over OFDM, the PAPR remains high with regular quadrature amplitude modulation (QAM), motivating the need for further low-PAPR techniques in both research literature and 3GPP standardization. One approach is to modify the symbol constellation, such as employing $\pi/2$-BPSK (supported in 5G uplink in addition to legacy 4G constellations), or other rotated-QAM schemes [4, 5]. Another add-on feature is to apply a spectral window, referred to as frequency-domain spectral (FDSS), reducing subsequently the PAPR at the cost of controlled self-interference. This is implicitly supported in 5G – for $\pi/2$-BPSK from its first release via relaxed spectrum error vector magnitude (EVM) mask requirements for proprietary signal implementations [6], and recently extended to QPSK in 5G-Advanced Rel. 18 by allowing power boosting – enabling FDSS as a transparent PAPR-reduction method without specifying an exact window.
Spectrum extension (SE) is another complementary PAPR-reduction technique, in which the DFT-spread data sequence is periodically extended – typically before applying an FDSS window. SE has been discussed in 4G [7, 8], considered for 5G [9, 10], and now proposed again for 6G [11, 12, 13]. The motivation is to enable low-PAPR also for higher-order modulation such as QPSK (i.e. 4-QAM), thereby improving coverage at higher-data rates than those supported by $\pi/2$-BPSK currently in 5G.
FDSS with SE (FDSS-SE) has been investigated in several works, e.g., [8, 14, 4, 10]. Early studies [7, 8, 14], often coupled the FDSS window design to the SE size, implicitly assuming a fixed SE size without adaptation to UE capability or other transmission parameters. These works did not examine the effect of varying SE size for a given FDSS roll-off factor. Often, academic works with a more systematic approach than 3GPP standardization, such as [4], optimize the FDSS window for a given system configuration, which does not align well with practical flexibility of current systems. Moreover, [7, 8] assumed tight spectrum-flatness requirements, restricting FDSS transition to narrow bands – limiting potential benefits compared to the broader transition bands allowed by the looser spectrum masks introduced in 5G. SE was also implicitly assumed to reduce PAPR at the expense of spectral efficiency, and so in order to mitigate this, preferably having extended subcarriers outside of the allocated band and overlapping with other transmissions.
The more recent work [10], which had a strong influence on the 5G Rel-18 study on FDSS-SE for uplink coverage enhancement, evaluated FDSS-SE at a fixed spectral efficiency, i.e. compensating the bandwidth loss by using higher-rate channel codes, and using a double-side symmetric SE with a fixed SE size at 25% of the bandwidth. QPSK with FDSS-SE was shown to outperform QPSK with FDSS but without SE, meaning that the power back-off reduction gain outweighed the bandwidth efficiency loss. In that setup, extended subcarriers did not overlap with other bands, enabling frequency-diversity combining of repeated symbols at the receiver.
While earlier academic works such as [14] considered single-side extension, symmetric SE as in [10] – already proposed in early 3G standardization [7, 8] – has become the de facto approach in 3GPP [11, 12, 13]. Symmetric SE corresponds to applying a particular circular shift to the subcarrier coefficients before single-side SE. Alternatively, in 3GPP Rel.18 study, we proposed asymmetric SE with optimized cyclic shifts [15, 16], derived as a by-product of the optimized constellation rotation approach in [4, 5]. Recently, similar asymmetric SE approaches have been studied in [17, 18] and are proposed for 6G [19]. Cyclic shifting is easier to implement and maintains compatibility with legacy systems, while incurring only negligible precision loss compared to optimal constellation rotation [17].
Moreover, a common misconception is that larger SE always reduces the PAPR, with its size constrained only by bandwidth-efficiency concerns. In contrast, this paper draws on new insights into FDSS-SE design, showing that for a given FDSS window there exists both an optimal SE size for PAPR minimization – beyond which PAPR performance degrades, and an optimal SE size for rate maximization – below which rate decreases. Hence, if in 6G the FDSS windows remain proprietary and UE-specific as in 5G, i.e., implemented by manufacturers and not optimizable by the network, then enforcing a single SE size for all UEs may be suboptimal and even detrimental for some of them.
During 5G Rel. 18 standardization [15, 16], we already pointed-out that PAPR minimization could be achieved with different SE sizes depending on the considered FDSS window. However, evaluations in 3GPP were conducted only for a few (typically competing) designs with a limited and heterogeneous set of specific FDSS windows and configurations, so general performance trends and behavior were not well established. This paper refines the PAPR analysis from our earlier 3GPP contributions with stronger theoretical support, as well as a rate analysis showing that SE also helps mitigate the SNR loss from FDSS. We use a generic FDSS-SE model with arbitrary frequency-shifts, encompassing prior single-side and symmetric definitions, and consider two single-parameter FDSS window families to evaluate the performance behaviors as a function of FDSS attenuation. Namely, we introduce a single-parameter “deformed Hann” window family covering FDSS windows used in 3GPP; and we also use a Kaiser window proposed for DFT-s-OFDM in [20] as a well-known approximation of the prolate spheroidal wave functions for optimal time- and band- limited localization [21].
In passing, we revisit the relevance of the PAPR and the alternative cubic-metric (CM), the latter typically preferred in 3GPP RAN1 study but actually shown here to be non-monotonic with FDSS shaping and thus less likely to align with power-derating trends for FDSS-SE.
We then address the PAPR minimization for FDSS-SE with a fixed FDSS window, leveraging and adapting the bounds from [4, 5] to determine the optimal SE size and frequency shift. We show that symmetric extension in [10] is optimal for $\pi/2$-BPSK but not for regular QAMs, where single-side extension [14] performs better but still not optimal. The optimal shifts for regular QAMs are provided and its PAPR gain shown to be well-anticipated by the bounds. More importantly, we study the optimal SE size for PAPR minimization. This optimum depends strongly on the FDSS window, differs between $\pi/2$-BPSK and regular QAMs, and is largely independent of the SE’s frequency shift or the allocated bandwidth (in proportion). All together, the optimized shift and SE size enable us to evaluate the best PAPR gain of FDSS with SE compared to FDSS without SE, showing that FDSS-SE provides the best PAPR reduction with QPSK compared to $\pi/2$-BPSK or higher-order QAMs.
Next, we examine the spectral efficiency of FDSS-SE. While conventional wisdom would suggest that SE always reduces rate, we derive a generic capacity expression for DFT-s-OFDM with FDSS-SE and the combining receiver of [10]. Our generic FDSS-SE approach simplifies the prior formulation of this receiver, and the capacity result shows that there is another SE size for maximizing the rate, dependent of the FDSS window and independent of the bandwidth, but, unlike the PAPR-optimal SE size, dependent also of the SNR. The PAPR-optimal and rate-optimal SE sizes differ, but a balanced choice can improve both metrics simultaneously.
The rest of the paper is as follows: Section II presents the DFT-s-OFDM transmitter model with FDSS-SE. Section III discusses accurate PAPR computation, compare with the CM its relevance for FDSS-SE study. Section IV analyzes FDSS-SE optimization for PAPR minimization. Section V covers the receiver model and spectral efficiency aspects. The paper is concluded in Section VI.
II DFT-s-OFDM Transmission
II-A DFT-s-OFDM with FDSS and Spectrum Extension
A DFT-s-OFDM symbol with cyclic prefix (CP) is defined for samples $N_{\rm cp}\leq n\leq N_{\rm fft}-1$ as
$s[n]=\frac{1}{\sqrt{N_{\rm fft}}}\sum_{k=0}^{N_{\rm sc}-1}W[k]X^{\rm se}[k]e^{{\rm j}\frac{2\pi nk}{N_{\rm fft}}}$ (1)
where
$X^{\rm se}[k]=X[(k+L)\;{\rm mod}\;N_{\rm data}]$ (2)
$\text{for }k=0,\ldots,N_{\rm sc}-1,$
are the spectrum-extended version of the subcarrier coefficients
$X[h]=\frac{1}{\sqrt{N_{\rm data}}}\sum_{m=0}^{N_{\rm data}-1}x[m]e^{-{\rm j}\frac{2\pi hm}{N_{\rm data}}},$ (3)
$\text{for }h=0,\ldots,N_{\rm data}-1,$
which are the DFT precoding of constellation symbols $x[m]\in\mathcal{C}$, where
  • $N_{\rm data}$ is the number of modulation constellation symbols, which are assumed to be independently and identically distributed (i.i.d.) QAM symbols with zero-mean and unit average energy, i.e. $\mathbb{E}[|x[m]|^{2}]=1$,
  • $N_{\rm sc}$ is the number of modulated OFDM subcarriers,
  • $N_{\rm e}=N_{\rm sc}-N_{\rm data}$ is the number of subcarriers used for SE, referred as the SE size,
  • $L$ is a circular shift value with $({\rm mod}\;N)$ being the modulo-$N$ operator,
  • $W[k]$ is a FDSS window which is assumed real and symmetric, and normalized as [^1] Note that in [4], the FDSS window power is set to $N_{\rm data}$. The choice of normalization here makes the transmit power constant and independent of the SE size and only dependent of the bandwidth allocation. This matches practical consideration as considered in 3GPP study on coverage enhancement.
    $\sum_{k=0}^{N_{\rm sc}-1}|W[k]|^{2}=N_{\rm sc}$ (4)
  • $N_{\rm fft}$ is the IFFT size of the OFDM modulation and $N_{\rm cp}$ is the CP length.
The model encompasses different SE approaches. For example, in [10], the spectrum extension is defined as double-side and symmetric with $L=N_{\rm data}-N_{\rm e}/2=-N_{\rm e}/2\;({\rm mod}\;N_{\rm data})$, while in [14, 4] extension is one-side with $L=0$. This interpretation as symmetric/asymmetric or two-/single-side is to a large extent artificial, and can be applied to any shift value. Indeed, as illustrated in Fig. 1, for any $L$, all SE definitions have the same symmetries, and can be seen as one- or two-side extensions. The complete set of symbols $\{X[0],\ldots,X[N_{\rm data}-1]\}$ is always in the “in-band”, as defined in [10], up to a cyclic-shift by $L+N_{\rm e}/2$ symbols; and the left-side excess-band symbols are always the repetition of symbols of the right-side in-band edge, similarly for the right-side excess band. The benefit of symmetric SE as in [10] may lie only in a more convenient indexing.
Figure 1: Illustration of spectrum-extended data sequence as a function of the shift parameter $L$ with $N_{\rm data}=10$ symbols and $N_{\rm e}=4$.
II-B Pulse Time-Multiplexing Intepretation
The DFT-s-OFDM transmission can be reformulated as a time-multiplexing of $N_{\rm data}$-pulses $\{p_{m}[n]\}_{m=0}^{N_{\rm data}-1}$. By combining (3), (2) and (1), we obtain
$s[n]=\frac{1}{\sqrt{N_{\rm fft}}}\sum_{m=0}^{N_{\rm data}-1}x[m]p_{m}[n]$ (5)
with time-domain pulses
$p_{m}[n]=e^{-{\rm j}\frac{2\pi L}{N_{\rm data}}m}p_{0}\left[n-\frac{N_{\rm fft}}{N_{\rm data}}m\right]$ (6)
which are different time and phase shifted versions of the same kernel pulse-shaping filter
$p_{0}[n]=\frac{1}{\sqrt{N_{\rm data}}}\sum_{k=0}^{N_{\rm sc}-1}W[k]e^{{\rm j}\frac{2\pi k}{N_{\rm fft}}n}.$ (7)
In the case of no FDSS windowing, $W[k]=1$, $\forall k$, we recover the conventional DFT-s-OFDM pulse, in the form of a Dirichlet kernel, with $N_{\rm sc}$ modulated subcarriers:
$p_{0}[n]=\frac{e^{{\rm j}\frac{\pi}{N_{\rm fft}}n(N_{\rm sc}-1)}}{\sqrt{N_{\rm data}}}\frac{\sin\left(\pi\frac{N_{\rm sc}}{N_{\rm fft}}n\right)}{\sin\left(\frac{1}{N_{\rm fft}}n\right)}.$ (8)
and then the pulses (6) are orthogonal; otherwise, with FDSS, they are in general not orthogonal anymore.
II-C FDSS Windows
We consider FDSS window families with convenient single shaping parametrization that characterizes their maximum FDSS power ripple defined as $20\log_{10}\frac{\min_{k}W[k]}{\max_{k}W[k]}$ [dB].
Deformed Hann Window: We define a deformed Hann window with shaping parameter $0\leq\beta\leq 1$,
$W_{\beta}[k]=\frac{1}{\omega}\left(1-\frac{1-\beta}{1+\beta}\cos\left(\frac{2\pi k+\pi}{N_{\rm sc}}\right)\right)$ (9)
where the normalization factor is $\omega=\sqrt{1+\frac{(1-\beta)^{2}}{2(1+\beta)^{2}}}$.
With $\beta=0$, we recover the classical Hann window up to normalization, while with $\beta=1$ corresponds to no FDSS shaping. This window is derived based on 3-tap filters commonly-used by industry in 3GPP, see Appendix -A. In this convenient form, the maximum power ripple of this FDSS window is tunable and directly given by its shaping parameter as $\approx 20\log_{10}\beta$. Moreover, the FDSS windows from 3-tap filters are not exactly symmetric and we applied an half index shift to correct this, which has a noticeable effect only for very narrow bands. Specifically, the specific 3-tap filter used in [22, 23, 24] are equivalently obtained as $W_{\beta}[k-1/2]$ using (9) with $\beta=-11$ and 14-14 dB, respectively.
Kaiser Window: The Kaiser window proposed for DFT-s-OFDM in [20] is defined with shaping parameter $\kappa$ as
$W_{\kappa}[k]=\frac{1}{\omega I_{0}\left(\kappa\right)}I_{0}\left(\kappa\sqrt{1-\frac{(k-\gamma)^{2}}{\gamma^{2}}}\right)$ (10)
where $\gamma=\frac{N_{\rm sc}-1}{2}$ and $I_{0}(\cdot)$ represents the zeroth-order modified Bessel function of the first kind; and $\omega$ is the normalization coefficient. The maximum of the window is at $W_{\kappa}[N_{\rm sc}/2]\approx 1/\omega$, while its minimum is at $W_{\kappa}[0]=\frac{1}{\omega I_{0}\left(\kappa\right)}$, so that its maximum power ripple is $\approx-20\log_{10}\left(I_{0}\left(\kappa\right)\right)$.
This window family was considered by us in 3GPP Rel. 18 study [15]. It is often close in shape to the truncated RRC (TRRC) design proposed in [10], considered also in 3GPP Rel. 18. However, the TRRC window is slightly less convenient for study as it has two parameters; and for a given parameter set, a shape that changes with the bandwidth size when the bandwidth is small.
Fig. 2 shows the variation of the window shapes obtained by the Hann and Kaiser window types for different $\beta$ and $\kappa$, along the maximum FDSS attenuation permitted with $\pi/2$-BPSK according to the specification of the EVM equalizer spectral flatness requirements in [6].
(a) Deformed Hann window
(b) Kaiser window
Figure 2: Two FDSS window types with single shaping parameters characterizing their maximum FDSS power ripple.
III PAPR: Definition and Relevance
Given an OFDM symbol, we define its PAPR as
${\rm PAPR}=\frac{\max_{0\leq n<N_{\rm fft}}|s[n]|^{2}}{\mathbb{E}\left[|s[n]|^{2}\right]}.$ (11)
where $\mathbb{E}\left[\cdot\right]$ is the statistical expectation. In numerical evaluations, we will use $N_{\rm fft}=2048$. PAPR of the digital OFDM signal provides an accurate PAPR estimation of the corresponding transmitted analog OFDM signal when oversampling is at least $N_{\rm fft}/N_{\rm data}=4$ [25].
Before analyzing the PAPR of FDSS-SE, we first point out some considerations on the definition and relevance of the PAPR metric, which appear to be insufficiently examined in the literature.
III-A Statistical Mean Power versus Instantaneous Mean Power
The PAPR in (11) is computed using the statistical mean power (i.e., the expectation in the denominator) as in many academic works, e.g., [14, 25, 4]. Alternatively, several other works define the PAPR using the instantaneous arithmetic mean power over a considered signal lengh $N$, i.e. ${\rm PAPR}^{\prime}=\frac{{\max_{0\leq n<N}}|s[n]|^{2}}{\frac{1}{N}\sum_{n=0}^{N-1}{|s[n]|^{2}}}$, and for this, often the signal length is one OFDM symbol as, e.g., in [11].
With FDSS, the discrepancy between these two PAPR definitions becomes more pronounced, especially when the number of subcarriers and/or OFDM symbols is small. This is because higher-order constellations combined with FDSS require more statistical realizations of the constellation symbols to average out the inherent randomness.
While the instantaneous average power may be argued to be more meaningful since it characterizes the PAPR for a specific signal realization as experienced by hardware; practical OFDM transmissions rarely consist of a single OFDM symbol, and the overall signal length can vary. Therefore, for a generic and analytically consistent evaluation based on a single OFDM symbol, the statistical mean power definition is preferable, if possible.
Fig. 4 illustrates these considerations with $N_{\rm sc}=24$ and deformed Hann window, where the $10^{-3}$ complementary cumulative distribution function (CCDF) level of the PAPR is plotted along increasing FDSS shaping attenuation. The PAPR is computed for both a single OFDM symbol and a 14 CP-OFDM symbol signal [^2] The maximum in the numerator of (11) is applied on the corresponding signal, i.e. for 14 OFDM symbols the peak is taken over a larger sample set, and so its PAPR is larger than with one OFDM symbol.. In the absence of FDSS (0 dB), both PAPR definitions yield the same value for $\pi/2$-BPSK and QPSK. However, as FDSS shaping increases, the two PAPR definitions diverge. This difference is small for $\pi/2$-BPSK, and for QPSK when signals comprise 14 CP-OFDM symbols. Similarly, for a single OFDM symbol with a larger number of subcarriers, the PAPR based on instantaneous mean power would approach that based on statistical mean power. However, as shown, for the case of QPSK with FDSS, the PAPR computed using the instantaneous mean power is significantly higher for single narrowband OFDM symbol than other cases, and thus fails to consistently reflect performance trends.
(a) $\pi/2$-BPSK
(b) QPSK
Figure 3: Inconsistencies among PAPR definitions when considering FDSS and QPSK in narrowband.
Figure 4: Inconsistent behavior of CM and PAPR for FDSS-SE with QPSK.
III-B Linear Fitting of Power De-rating by PAPR and CM
The PAPR measures the envelope fluctuation of a signal. Peaks in signal envelope either cause distortion by entering the nonlinear region of the power amplifier; or alternatively require power de-rating to avoid such distortion which directly reduces the link SNR. 3GPP standards specify a maximum power reduction (MPR) level that limits the allowed de-rating (also referred to as output back-off or offset) to ensure conformance with RF requirements, including those for signal distortion and adjacent channel interference. In practical systems, the required power de-rating depends on the bandwidth allocation size and its position within the channel bandwidth, typically being higher for either large allocations or small allocations located near the channel edge. Thus, in the corresponding 3GPP specification [26], edge allocations have larger MPR values compared to inner allocations, and larger allocated bandwidth carry additional MPR margins. In general, higher modulation orders lead to higher PAPR and thus allow greater MPR values.
The cubic metric (CM) is an alternative to PAPR that has been considered superior in 3GPP for quantifying MPR. The (raw) cubic metric is defined as ${\rm CM}\,{\rm[dB]}=20\log_{10}{\rm CM}$ with ${\rm CM}={\rm rms}\left(\left|\frac{s[n]}{{\rm rms}(s[n])}\right|^{3}\right)$ where ${\rm rms}\left(s[n]\right)=\sqrt{\sum_{n=0}^{N-1}|s[n]|^{2}}$. The metric was introduced during late 3G UMTS standardization, based on nonlinear amplifier modeling, with the argument that it more accurately predicts power de-rating than PAPR [27]. From its inception, CM has been paired with linear regression coefficients to match MPR measurements. In fact, the 3G UMTS specification [28] includes multiple shifted, scaled, and quantized versions of CM depending on the system configuration. For 4G LTE, the CM fitting was updated [29] based on OFDM signal measurements, and different regression coefficients were recommended for two different bandwidth configurations. Since then, 5G NR has reused these definitions from [27], originally derived for 4G – and in some cases from 3G work, see e.g., [30]. However, applying the early-4G linear fit often leads to negative power de-rating estimates when evaluating modern low-PAPR waveforms, such as DFT-s-OFDM with $\pi/2$-BPSK and FDSS as well as other schemes see e.g. [31], which raises questions about the practical meaning of these CM-based MPR estimate.
It is worth emphasizing that CM may provide much better power de-rating estimation than PAPR, primarily because it is always used with fitted regression coefficients from measurements, whereas PAPR is typically evaluated in its raw form, without linear fitting. However, nothing prevents applying a linear regression to PAPR as well, to estimate more realistic de-rating values. In fact, in the original 3GPP contribution introducing CM [27], a linear fitting of PAPR was also considered: ${\rm MPR}\approx 0.855({\rm PAPR}-3)$. Although CM provided a slightly better overall fit, the relationship between power de-rating and PAPR still followed a linear correlation reasonably well. More recent evaluations, such as [32], confirms this.
In any case, neither CM nor PAPR can fully predict realistic power de-rating values. For instance, in [24], both the CM and PAPR of QPSK with FDSS appear worse than with clipping, while the measured power de-rating is actually better. Most importantly, the impact of bandwidth allocation size and position relative to the system bandwidth is not captured by either CM or PAPR.
III-C Non-Monotonic CM Behavior with FDSS
In 3GPP Rel-18 study on coverage enhancement, CM-based power de-rating estimation was considered by RAN1 as an intermediate step prior to RAN4 MPR evaluations [15]. This CM-based approach initially influenced the FDSS-SE designs in RAN1, but its reliability was later challenged, and many RAN1 participants eventually relied on direct MPR evaluations [16, 24].
We point-out a malfunctioning behavior of CM when applied with FDSS and higher-order QAM modulations: as FDSS shaping increases, CM follows a bell-shaped curve. As shown in Fig. 4, unlike PAPR, which improves systematically with larger FDSS attenuation, CM exhibits this trend only for $\pi/2$-BPSK, but not for QPSK. This inconsistency between PAPR and CM is further observed in Fig. 4 for FDSS-SE and a larger bandwidth ($N_{\rm sc}=96$, $N_{\rm e}=10$), where the bell-shape of CM becomes even more pronounced.
This behavior does not align with an expected power de-rating prediction. Intuitively, as the signal spectrum becomes more confined, adjacent channel interference decreases, which is one of the main conformance issues limiting the transmit power. In fact, evaluations from 3GPP Rel-18 study [24, 16] confirms that power de-rating gains are larger with more aggressive windows. Moreover, it is unlikely that a more confined spectrum would lead to power de-rating gain with $\pi/2$-BPSK but a loss with QPSK.
Realistic MPR evaluations are beyond the scope of this paper, but this suggests that the presumed superiority of CM over PAPR for power de-rating prediction should be reconsidered in the context of FDSS-SE. Prior works such as [32], which showed better prediction with CM or other metrics, did not account for the impact of FDSS. Therefore, in this paper, we retain the classical PAPR metric as, not necessarily the most accurate, but a consistent and analytically tractable indicator of power de-rating performance trends.
IV Spectrum Extension Minimizing the PAPR
In this section, we study the PAPR of FDSS-SE as a function of the SE size $N_{\rm e}=N_{\rm sc}-N_{\rm data}$ and the frequency-shift $L$, assuming that $N_{\rm sc}$ and the FDSS window are pre-configured and fixed.
IV-A PAPR Bounds
We begin by expressing PAPR bounds that will help to characterize and verify the PAPR behavior. Based on the pulse multiplexing interpretation (5) with pulses $p_{m}[n]$ defined in (6), a simple adaptation and generalization of the approach in [4] and [5] (see Appendix -B) yields the following PAPR upper bound, which depends only on $N_{\rm e}$, $L$, and the symbol constellation.
Lemma 1.
For an SE size $N_{\rm e}$ with shift $L$ and constellation $\mathcal{C}$ with largest amplitude $A_{\mathcal{C}}$, we have
${\rm PAPR}\leq{\rm PAPR}^{\rm U}(\mathcal{C},N_{\rm e},L)\leq{\rm PAPR}^{\rm GU}(A_{\mathcal{C}},N_{\rm e})$ (12)
with
${\rm PAPR}^{\rm U}(\mathcal{C},N_{\rm e},L)=\frac{A_{\mathcal{C}}^{2}}{N_{\rm sc}}\max_{n}\left\{\sum_{i,j=0}^{N_{\rm data}-1}|p_{i}[n]||p_{j}[n]|u_{i,j}^{\mathcal{C}}\right\}$ (13)
where $p_{i}[n]$ are the time-domain pulses (6) and
$u_{i,j}^{\mathcal{C}}=\max_{\omega\in\Omega_{\mathcal{C}}}\left|\cos\left((i-j)\left[\phi-\frac{(2L+N_{\rm e}-1)}{N_{\rm data}}\pi\right]+\omega\right)\right|,$ (14)
$\Omega^{\mathcal{C}}$ being the set of phase differences among constellation symbol modulo $\pi$.
This is further upper-bounded by the more general bound
${\rm PAPR}^{\rm GU}(A_{\mathcal{C}},N_{\rm e})=\frac{A_{\mathcal{C}}^{2}}{N_{\rm sc}}\max_{n}\left\{\sum_{i,j=0}^{N_{\rm data}-1}|p_{i}[n]||p_{j}[n]|\right\}.$ (15)
More precisely:
  • For $\pi/2$-BPSK, $\phi=\frac{\pi}{2}$, $A_{\mathcal{C}}=1$, and $\Omega_{\mathcal{C}}=\{0\}$.
  • For QPSK, $\phi=0$, $A_{\mathcal{C}}=1$, and $\Omega^{\mathcal{C}}=\{0,\frac{\pi}{2}\}$.
  • For 16-QAM, $\phi=0$, $A_{\mathcal{C}}^{2}=1.8$, and $\Omega^{\mathcal{C}}$ contains 10 different values.
As the constellation grows, the set of angles $\Omega^{\mathcal{C}}$ becomes a denser sampling of $0,\,\pi]$ so the first upper bound converges to the general upper bound ${\rm PAPR}^{\rm U}(\mathcal{C},N_{\rm e},L)\to{\rm PAPR}^{\rm GU}(A_{\mathcal{C}},N_{\rm e})$.
Moreover, for regular QAM, note that the maximum symbol power $A_{\mathcal{C}}^{2}$ occurs only for a subset of constellation points at the four edges , forming a QPSK sub-constellation with larger amplitude. We can therefore expect that symbol combination of the this sub-QPSK constellation will yield some of the highest signal peaks, leading to:
${\rm PAPR}^{\rm U}(\mathcal{C}_{\rm QAM},N_{\rm e},L)\approx A_{\mathcal{C}_{\rm QAM}}^{2}\times{\rm PAPR}^{\rm U}(\mathcal{C}_{\rm QPSK},N_{\rm e},L)$ (16)
To compute these bounds, the maximum over one OFDM symbol duration $0\leq n\leq N_{\rm fft}-1$ must be found. The computation complexity of this step can be greatly reduced by noting that the function inside the max-operator is periodic, since all pulses in (6) are regular time-shifted versions of the same kernel. It is therefore sufficient to search over $N_{\rm fft}/N_{\rm data}$ consecutive samples to find the maximum.
IV-B Best Shifting Value
The best shifting value $L$ for FDSS-SE follows from the phase difference between two neighboring pulses with indices $m$ and $(m+1)$ (see Appendix -C):
$\angle\frac{p_{m+1}[n]}{p_{m}[n]}=-\frac{2\pi}{N_{\rm data}}\left(L+\frac{(N_{\rm e}-1)}{2}\right)\;({\rm mod}\,\pi),$ (17)
which corresponds to the term (14) inside the bound (13).
For $\pi/2$-BPSK, the purpose of the $\pi/2$-rotation is to ensure that BPSK symbol of neighboring pulse are transmitted with a phase difference of approximately $\frac{\pi}{2}\;({\rm mod}\,\pi)$, such that their maximum possible power combining is minimized. Therefore, the best $L$ in this case is found such that it does not have any effect on the phases and (17) is as close to zero as possible, i.e.
$L$ $=$ $\left\lfloor\frac{\lambda}{2}N_{\rm data}-\frac{N_{\rm e}-1}{2}\right\rceil$ (18)
$=$ $\left\lfloor\frac{\lambda}{2}N_{\rm sc}-\frac{N_{\rm e}-\lambda-1}{2}\right\rceil.$
where $\lambda$ is an integer and $\left\lfloor\cdot\right\rceil$ returns the closest integer. With $\lambda=2$ and $N_{\rm e}$ even, then one can select $L=N_{\rm data}-\frac{N_{\rm e}}{2}$ corresponding to symmetrical double-side SE considered in [10].
For QPSK and other QAMs, choosing $L$ as in [10] is, however, sub-optimal, and most other $L$ values can provide slightly better PAPR. In these cases, the value of $L$ providing the lowest PAPR depends on the spectrum extension size $N_{\rm e}$ but also on the total bandwidth size via $N_{\rm data}$ or $N_{\rm sc}$. Here, the lowest PAPR is obtained by creating a phase difference between neighboring pulses close to $\pi/4\;({\rm mod}\,\pi/2)$. Namely, one should have $\frac{2\pi}{N_{\rm data}}\left(L+\frac{(N_{\rm e}-1)}{2}\right)\approx\frac{\pi}{4}\;\left({\rm mod}\,\frac{\pi}{2}\right).$ Therefore, the lowest PAPR should be achieved by selecting $L$ satisfying
$L$ $=$ $\left\lfloor\frac{(2\lambda+1)}{8}N_{\rm data}-\frac{N_{\rm e}-1}{2}\right\rceil$ (19)
$=$ $\left\lfloor\frac{(2\lambda+1)}{8}N_{\rm sc}-\frac{(2\lambda+5)}{8}N_{\rm e}+\frac{1}{2}\right\rceil$
where $\lambda$ is an integer.
Figure 5: Example of PAPR fluctuation as a function of different frequency-shift value $L$ for $\pi/2$-BPSK, QPSK and 16-QAM.
Figure 6: PAPR deviation of QPSK for different $L$ as a function of SE ratio $N_{\rm e}/N_{\rm sc}$.
Fig. 5 shows the $10^{-3}$-CCDF PAPR for $\pi/2$-BPSK, QPSK, and 16-QAM as a function of $L$ with $N_{\rm sc}=96$, $N_{\rm e}=10$. The FDSS is the deformed Hann window with $\beta=-11$ dB. The bounds (13) for $\pi/2$-BPSK and QPSK are also shown. The PAPR curves and their bounds have several equivalent minima as predicted by (18) for $\pi/2$-BPSK and by (19) for both QPSK and 16-QAM. Note that for $\pi/2$-BPSK, the decimal part of (18) is $0.5$ so rounding up or down to the nearest integer is valid. One can observe that the choice of $L$ can have a large impact on the PAPR for $\pi/2$-BPSK, but this impact is more moderate for higher-order QAM.
In Fig. 6, we show the deviation of the $10^{-3}$-CCDF PAPR as a function of the SE ratio $N_{\rm e}/N_{\rm sc}$ for QPSK with $N_{\rm sc}=96$ and Kaiser FDSS window with $\kappa=2$, which corresponds to a maximum power attenuation of $7.15$ dB. The PAPR deviation between i) $L$ from (18) corresponding to the symmetric extension in [10], and ii) $L$ from (19), is computed with reference to iii) $L=0$, corresponding to the single-side extension in [14]. The figure also show a prediction of this PAPR deviation using the bound (13) with the corresponding $L$ values. As seen, for QPSK the symmetric extension from [10] is always worse than the single-side extension from [14]. Similar results can be obtained for higher-order QAMs. The maximum PAPR gain over [10] by using a different $L$ is only a fraction of a dB, here up to 0.5 dB, and this gain decreases as the SE size increases.
IV-C Optimum SE Size $N_{\rm e}^{({\rm min\,PAPR})}$ for Minimum PAPR
IV-C1 Existence of an Optimal SE
It is well known in the literature that SE can further reduce the PAPR on top of FDSS. However, it less recognized that there exists an optimum SE size that minimizes the PAPR, and that a very large SE may even increase it.
To intuitively understand this, consider the pulse-shaping filters in (6), derived from the band-limited FDSS window $W[k]$ in (7). Their amplitude essentially follows a sinc shape, with more or less attenuated side lobes. The kernel filter shape in (7) is independent of $N_{\rm e}$, except for a scaling factor $\frac{1}{N_{\rm sc}-N_{\rm e}}$, which causes the peak energy of each pulse to increase as $N_{\rm e}$ increases. In parallel, since there are $(N_{\rm sc}-N_{\rm e})$ pulses per OFDM symbol, the pulses are spaced $\frac{N_{\rm fft}}{N_{\rm sc}-N_{\rm e}}$ samples apart in (6). This spacing increases with $N_{\rm e}$, which in turn changes the degree of pulse overlap due the sinusoidal shape of the pulses. Both effects combined, depending of $N_{\rm e}$, their overall combining may decrease or increase as a function of $N_{\rm e}$, leading to a non-monotonic PAPR variation.
In Fig. 7, we illustrate this with a deformed Hann FDSS window ($\beta=-11$ dB and $N_{\rm sc}=48$). Three groups of three pulses are shown for different $N_{\rm e}$, along with their respective non-coherent combining. Increasing the SE size to $N_{\rm e}=12$ raises the individual pulse energy but also spreads the pulses farther apart, reducing overlap and lowering their combined peak. However, further increasing to $N_{\rm e}=28$ causes the larger individual pulse energy to dominate, raising the combined peak beyond the level without SE ($N_{\rm e}=0$).
Figure 7: Groups of 3 pulses and their non-coherent combining for different values of $N_{\rm e}$.
(a) $N_{\rm sc}=96$, No FDSS, QPSK
(b) $N_{\rm sc}=96$, Kaiser $\kappa=2$, QPSK
(c) $N_{\rm sc}=96$, Hann $\beta=-11$ dB, QPSK
(d) $N_{\rm sc}=96$, Kaiser $\kappa=2$, $\pi/2$-BPSK
(e) $N_{\rm sc}=96$, Kaiser $\kappa=2$, 16-QAM
(f) $N_{\rm sc}=48$, Kaiser $\kappa=2$, QPSK
Figure 8: Examples of PAPR as a function of SE $N_{\rm e}$ showing the eixtence of a non-trivial SE size $N_{\rm e}^{({\rm min\,PAPR})}$ for PAPR minimization.
In Fig. 8, we plot several examples of PAPR versus the SE ratio $N_{\rm e}/N_{\rm sc}$. Each plot shows the maximum PAPR from $10^{6}$ Monte Carlo points, the $10^{-1}$, $10^{-2}$, and $10^{-3}$ CCDF, and the bounds from (13) and (15). All subplots is with QPSK, except Fig. 8(d) with $\pi/2$-BPSK and Fig. 8(e) with 16QAM. All have $N_{\rm sc}=96$ except Fig. 8(f) with $N_{\rm sc}=48$. The first row, i.e. Figs. 8(a)–(c), uses different FDSS windows, while the second row, i.e. Figs. 8(d)–(f), varies the modulation or bandwidth, all with a Kaiser FDSS window ($\kappa=2$). Shift values match single-side extension ($L=0$) in the literature, except Fig. 8(d) where $L$ is chosen according to (18).
Fig. 8 clearly shows that the PAPR does not decrease monotonically with $N_{\rm e}$. In all cases, the PAPR decreases up to an optimum SE size $N_{\rm e}^{({\rm min\,PAPR})}$, beyond which it increases again. Selecting $N_{\rm e}>N_{\rm e}^{({\rm min\,PAPR})}$ is detrimental not only for spectral efficiency but also for PAPR reduction. We have systematically verified the existence of this global optimum across many scenarios. Remark that even without FDSS [Fig. 8(a)], SE can provide some PAPR reduction and so $N_{\rm e}^{({\rm min\,PAPR})}>0$.
From Figs. 8(a)–(c), one can observe that $N_{\rm e}^{({\rm min\,PAPR})}$ changes with the FDSS window, while the PAPR curves are almost identical between QPSK [Fig. 8(b)] and 16-QAM [Fig. 8(e)]. Similarly, the PAPR behavior is consistent for QPSK with different bandwidths [Figs. 8(b) and 8(f)]. In [15], we showed by tedious simulations that $N_{\rm e}^{({\rm min\,PAPR})}$ as a percentage of $N_{\rm sc}$, remains roughly constant for different $N_{\rm sc}$, given different examples of FDSS windows. In the following, we show that PAPR bounds can provide a good characterization of this optimum without exhaustive simulations.
IV-C2 Characterizing $N_{\rm e}^{({\rm min\,PAPR})}$
As observed in Fig. 8, the PAPR behavior as a function of $N_{\rm e}$ is well characterized by the bounds (13) and (15). Accordingly, the value of $N_{\rm e}^{({\rm min\,PAPR})}$ can be closely approximated using bound (13), with slightly better precision than bound (15). For Fig. 8(e) with 16QAM, we also display the approximate bound (16) which appears to be a slightly better and simpler way to approximate $N_{\rm e}^{({\rm min\,PAPR})}$ than (13) in this case. Note, however, that the global optimum SE size obtained from simulations may correspond only to a local minimium in the PAPR bound, as shown in Fig. 8(a) – something we observed for smaller bandwidths and FDSS windows with small attenuation. A simple way to circumvent this issue is to restrict the search range for small FDSS attenuation, or inject a correction factor into the bound. Indeed, we observe that the bounds are further apart for $N_{\rm e}=0$ and become tighter as $N_{\rm e}\to N_{\rm sc}$. A practical method to compensate for this discrepancy is to approximate the PAPR as ${\rm PAPR}\;{\rm[dB]}\approx{\rm PAPR}^{\rm U}(\mathcal{C},N_{\rm e},L)\;{\rm[dB]}-K\left(1-\frac{N_{\rm e}}{N_{\rm sc}}\right)$ where $K$ is the gap between the bound and simulation at $N_{\rm e}=0$, which can empirically be evaluated.
(a) Different QAM order or bandwidth
(b) Different $L$ for QPSK and $\pi/2$-BPSK
Figure 9: Optimum SE size $N_{\rm e}^{({\rm min\,PAPR})}$ for different FDSS windows.
[…]
In Fig 9(b), we show that while the shift value of $L$ can impact the PAPR, the optimal SE size is nearly the same for different $L$ values. For QPSK, $N_{\rm e}^{({\rm min\,PAPR})}$ with $L$ as in (18), (19) and $L=0$ is obtained from simulations, together with the corresponding bound-based estimates. All cases are closely overlapping, and thus the bound (13) does not capture well the small variation of $N_{\rm e}^{({\rm min\,PAPR})}$ with different $L$ values. In Fig 9(b), $N_{\rm e}^{({\rm min\,PAPR})}$ is also shown for $\pi/2$-BPSK with $L$ as in (18) and $L=0$, from simulation and corresponding bounds. In this case, the range of $N_{\rm e}^{({\rm min\,PAPR})}$ is much smaller and almost constant around $5\%$, indicating that only a small amount of SE may be beneficial. Moreover, contrary to QPSK and other QAMs, as FDSS shaping increases, $N_{\rm e}^{({\rm min\,PAPR})}$ decreases.
IV-C3 Maximum PAPR gain of FDSS-SE over FDSS
Finally, Fig. 10 shows the $10^{-3}$ PAPR CCDF as a function of FDSS shaping, both with and without SE. To reach the lowest PAPR values, the shift parameter $L$ is according to (18) for $\pi/2$-BPSK and (19) for QPSK and 16-QAM; and for FDSS-SE, the corresponding optimum SE sizes $N_{\rm e}^{({\rm min\,PAPR})}$ estimated by simulation are used. For QPSK, both filter families – Hann and Kaiser windows – are considered, while for $\pi/2$-BPSK and 16-QAM only results with Hann window are shown for clarity.
For QPSK, SE with Hann window provides an almost constant PAPR gain over FDSS without SE from $1.74$ dB to $1.54$ dB. With Kaiser window, the gain of SE is the same with no FDSS but decreases slightly more to $1.31$ dB. For both window types, the lowest PAPR achieved with FDSS-SE is approximately the same for a given maximum window attenuation ripple. For 16-QAM, the PAPR gain with Hann window is also constant but smaller, at about 1.31.3dB. For $\pi/2$-BSPK, the behavior is different: the gain of SE over FDSS gradually decreases with larger shaping, from 1.51.5dB to 0.30.3dB.
This is consistent with the consensus during 3GPP 5G studies that SE on top of FDSS provides little benefit for $\pi/2$-BSPK [16], especially since in 5G the FDSS windows have been allowed to have such levels of attenuation ripple. Overall FDSS-SE provides the greatest PAPR reduction potential for QPSK, reaching values as low as 2-3 dB – similar to those obtained with $\pi/2$-BSPK using FDSS alone.
Figure 10: Comparisons of the minimum PAPR achieved with and without SE over different FDSS attenuation.
V Spectrum Extension Maximizing the Transmission Rate
The PAPR reduction from FDSS comes at the cost of self-induced inter-carrier interference (ICI) as the time-multiplex pulses are not orthogonal anymore, which reduces the transmission rate. The PAPR reduction obtained by SE comes at the cost of a smaller effective bandwidth, which can also degrade the transmission rate – although, as we will show, not always. In this section, we discuss the achievable rate of FDSS-SE as a function of the SE size.
V-A DFT-s-OFDM Reception
The transmitted signal $s[n]$ is assumed to be received over a $P$-path channel $\{h_{p}\}_{p=0}^{P-1}$, with $P<N_{\rm cp}$, such that the received OFDM symbol for $0\leq n\leq(N_{\rm fft}-1)$ is
$y[n]=\sqrt{\mathsf{snr}}\sum_{l=0}^{P-1}{h_{p}s[n-p]}+z[n].$ (20)
where $\mathsf{snr}$ denotes the signal-to-noise ratio (SNR), $\{z[n]\}$ is an average white Gaussian noise (AWGN) with $\mathbb{E}\left[|z[n]|^{2}\right]=1$, and the channel is normalized so that $\sum_{l=0}^{P-1}\mathbb{E}\left[|h_{p}|^{2}\right]=1$.
After FFT demodulation, the frequency-domain symbol on subcarrier $0\leq k\leq N_{\rm sc}-1$ is
$Y[k]$ $=$ $\frac{1}{\sqrt{N_{\rm fft}}}\sum_{n=0}^{N_{\rm fft}-1}y[n]e^{-{\rm j}2\pi\frac{kn}{N_{\rm fft}}}$ (21)
$=$ $H[k]X^{\rm se}[k]+Z[k]$
with per-subcarrier channel coefficient
$H[k]=\sqrt{\mathsf{snr}}W[k]\bar{H}[k]$ (22)
where $\bar{H}[k]=\sum_{p=0}^{P-1}h_{p}e^{-{\rm j}2\pi\frac{kp}{N_{\rm fft}}}$ is the DFT of the channel impulse response and $Z[k]=\frac{1}{\sqrt{N_{\rm fft}}}\sum_{n=0}^{N_{\rm fft}-1}z[n]e^{-{\rm j}2\pi\frac{kn}{N_{\rm fft}}}$ is the frequency-domain noise with unit variance.
V-A1 Equalization
We adopt the mixed MRC-MMSE combining from [10] which was the primary approach [^4] An alternative could be applying combining before equalization as $Y[k]+Y[k+N_{\rm data}]=(H[k]+H[k+N_{\rm data}])X^{\rm se}[k]+Z[k]+Z[k+N_{\rm data}]$ but this is worse as $|H[k]+H[k+N_{\rm data}]|^{2}\leq|H[k]|^{2}+|H[k+N_{\rm data}]|^{2}$. considered in the 3GPP Rel-18 study on uplink coverage enhancement, generalized here to any shift value $L$. With perfect channel knowledge, the symbols are first equalized via MRC:
$R[k]=H[k]^{*}Y[k]\quad\text{ for }0\leq k<N_{\rm sc}$ (23)
and then repeated symbols are combined as
$\tilde{R}[k]=\begin{cases}R[k]+R[k+N_{\rm data}]&0\leq k<N_{\rm e}\\ R[k]&N_{\rm e}\leq k<N_{\rm data}\end{cases}.$ (24)
The combining in (24) is equivalent to [10, Eq. (8)], but here it is simplified based on the symbol symmetry shown in Fig. 1, avoiding the need to define five sub-bands as in [10].
By substitution,
$R[k]=|H[k]|^{2}X^{\rm se}[k]+H[k]^{*}Z[k]$ (25)
and using the periodicity $X^{\rm se}[k+N_{\rm data}]=X^{\rm se}[k]$, we obtain
$\tilde{R}[k]$ $=$ $G[k]X^{\rm se}[k]+\tilde{Z}[k]\quad\quad\quad\quad\text{ for }0\leq k<N_{\rm data}$ (26)
$=$ $G[k]X[(k+L)\;{\rm mod}\,N_{\rm data}]+\tilde{Z}[k]$
where
$G[k]=\begin{cases}|H[k]|^{2}+|H[k+N_{\rm data}]|^{2}&0\leq k<N_{\rm e}\\ |H[k]|^{2}&N_{\rm e}\leq k<N_{\rm data}\end{cases}.$ (27)
and the equalized noise is
$\tilde{Z}[k]=\\ \begin{cases}H[k]^{*}Z[k]+H[k+N_{\rm data}]^{*}Z[k+N_{\rm data}]\;0\leq k<N_{\rm e}\\ H[k]^{*}Z[k]\quad\quad\quad\quad\quad\quad\quad\quad N_{\rm e}\leq k<N_{\rm data}\end{cases}.$ (28)
Accordingly, the MMSE equalization of the combined symbols is
$\tilde{R}_{\rm eq}[k]=\frac{\tilde{R}[k]}{G[k]+1}.$ (29)
V-A2 DFT Despreading
The frequency shift is then reverted as [^5] Since this cyclic shift is equivalent to a constellation phase rotation, an alternative is to first apply the IDFT, followed by linear phase compensation on the demodulated symbols.
$\tilde{R}_{\rm eq}^{\prime}[k]=\tilde{R}_{\rm eq}[k-L\;({\rm mod}\;N_{\rm data})]$ (30)
followed by the IDFT
$r[m]=\frac{1}{\sqrt{N_{\rm data}}}\sum_{k=0}^{N_{\rm data}-1}\tilde{R}_{\rm eq}^{\prime}[k]e^{{\rm j}\frac{2\pi}{N_{\rm data}}km}.$ (31)
V-B Achievable Rate with FDSS-SE
From Appendix -D, FDSS-SE yields the following achievable rate, that we will refer with a slight abuse of terminology as the FDSS-SE capacity.
Lemma 2.
The effective SINR of the demodulated symbol in (31) is independent of $m$ and given by
$\mathsf{SINR}^{\rm eff}=\frac{g_{0}}{1-g_{0}}$ (32)
where $g_{0}=\frac{1}{N_{\rm data}}\sum_{k=0}^{N_{\rm data}-1}\frac{G[k]}{G[k]+1}$ with combined channel gains $G[k]$ in (27), and the achievable rate is
$C=\frac{N_{\rm data}}{N_{\rm sc}}\mathbb{E}\left[\log_{2}\left(\frac{1}{1-g_{0}}\right)\right]\quad{\rm[bpcu]}.$ (33)
Figure 11: Simulated BER of QPSK versus theoterical BER using $\mathsf{SINR}^{\rm eff}$ in Lem. 2.
With the intention of verifying Lem. 2, Fig. 11 compares simulated bit error rates (BER) of Gray-mapped QPSK with theoretical BER of QPSK under $\mathsf{SINR}^{\rm eff}$ from (32), i.e., $Q\left(\sqrt{\mathsf{SINR}^{\rm eff}}\right)$ averaged over the different channel realizations. Channel assumptions are $N_{\rm sc}=96$, 3GPP TDL-C channel with $300$ ns delay spread and Rayleigh block fading. The near-perfect match confirms Lem. 2, indicating that the ICI behaves nearly Gaussian.
Remark 1.
For a basic receiver without SE symbol combining as discussed in [10], Lem. 2 still applies but with $G[k]=|H[N_{\rm e}/2+k]|^{2}$ for $0\leq k<N_{\rm data}$ instead of (27).
(a) RRC FDSS
(b) Hann FDSS
Figure 12: SNR loss of FDSS-SE wrt plain DFT-s-OFDM, at different spectral efficiency in [bpcu], estimated either from the capacity formula in Lem. 2 or from the BLER simulation in [15], as a function of the SE ratio $\frac{N_{\rm e}}{N_{\rm sc}}$.
V-C Optimum SE Size $N_{\rm e}^{({\rm max\,Capa})}$ for Maximum Rate
V-C1 Existence of an Optimal SE
The FDSS-SE capacity (33) has a pre-logarithm scaling factor that decreases with SE size. However, the term $g_{0}$ inside the logarithm may either decrease (due to a smaller summation range) or increase (due to larger $G[k]$ values) as a function of $N_{\rm e}$. Thus, as in the PAPR minimization case, a non-obvious optimum SE size $N_{\rm e}^{({\rm max\,Capa})}$ exists for capacity maximizing; and a too small SE size with $N_{\rm e}<N_{\rm e}^{({\rm max\,Capa})}$ may not achieve the maximum rate.
A close examination of 3GPP BLER results confirms that FDSS-SE performance is non-monotonic with SE size. For example, rate improvements with SE can be observed in [15, Tabs. 2-7], and the achievable-rate formulation in (33) can help to predict such performance behavior in realistic setups. Fig. 12 shows the SNR loss of FDSS-SE relative to plain DFT-s-OFDM versus the SE ratio for $N_{\rm sc}=96$ with 15 kHz subcarrier spacing, and a TDL-C(300 ns) channel. Two FDSS window types are considered: a 3-tap filter equivalent (up to half a subcarrier shift) to a deformed Hann window with $\beta=-11$ dB, and the TRRC window from [10] with parameters $(\rho,\beta)=(0.5,-0.65)$. The theoretical SNR loss is obtained by numerically finding the required SNR – with and without FDSS-SE – in (33) to achieve target spectral efficiencies of $1/3$, $2/3$, and $1$ bpcu. This is compared with 3GPP evaluations of the SNR loss reported in [15, Tabs. 2-7] for reaching $10^{-1}$ BLER with QPSK and code rates $1/6$, $1/3$, and $1/2$. We note that these BLER simulations are not from the present authors, and follows the practical 3GPP assumptions in [15, Tab. 1]. As shown in Fig. 12, the SNR loss also forms a bell-shaped curve as versus the SE ratio, yielding an optimal SE size $N_{\rm e}^{({\rm max\,Capa})}$ from rate-maximization perspective. The theoretical SNR-loss curves exhibit trends consistent with the 3GPP evaluations in [15] – particularly for half-rate coding (1 bpcu) – and can be used to characterize this optimum SE size.
Figure 13: Optimal SE size $N_{\rm e}^{({\rm max\,Capa})}$ in % of $N_{\rm sc}$ for two channel models and two bandwidth size, as function of SNR, given a deformed Hann FDSS window with $\beta=-11$ dB.
Figure 14: Comparison of achievable rate between DFT-s-OFDM with optimized FDSS-SE and with FDSS only, as a function of FDSS shaping.
(a) Optimum SE sizes, deformed Hann window
(b) Optimum SE sizes, Kaiser window
(c) PAPR, deformed Hann window
(d) PAPR, Kaiser window
Figure 15: Comparison of $N_{\rm e}^{({\rm max\,Capa})}$ and $N_{\rm e}^{({\rm min\,PAPR})}$, and the resulting PAPR trade-off as a function of FDSS shaping.
V-C2 Characterizing $N_{\rm e}^{({\rm max\,Capa})}$
Since the combined channel gains $G[k]$ directly depend on the FDSS, the optimal SE size for rate maximization depends on the FDSS shaping. In addition, as $\mathsf{snr}\to\infty$, $g_{0}\approx(1-f(N_{\rm e})/\mathsf{snr})$ for some bounded function $f(N_{\rm e})$ such that the impact of $N_{\rm e}$ on $g_{0}$ diminishes, and $C\approx\left(1-\frac{N_{\rm e}}{N_{\rm sc}}\right)\log_{2}(\mathsf{snr}/f(N_{\rm e}))$ such that $N_{\rm e}^{({\rm max\,Capa})}\to 0$ as $\mathsf{snr}\to\infty$ in order to maximize the pre-log factor of the capacity. Fig. 13 illustrates $N_{\rm e}^{({\rm max\,Capa})}$ (in % of $N_{\rm sc}$) versus SNR for a Hann FDSS window with $\beta=-11$ dB under both TDL-C (300 ns) delay spread and AWGN channels, with $N_{\rm sc}=96$ and $N_{\rm sc}=600$. The optimum SE size decreases with SNR and tends to zero. Similar to the PAPR case, the optimum ratio $N_{\rm e}^{({\rm max\,Capa})}/N_{\rm sc}$ is nearly bandwidth-invariant. However, under frequency-selective channel, it slightly decreases with larger bandwidth. Overall, $N_{\rm e}^{({\rm max\,Capa})}$ depends primarily on SNR and FDSS shaping.
V-C3 Maximum Rate of FDSS-SE over FDSS
Fig. 14 shows the maximum achievable rate of FDSS-SE when using $N_{\rm e}^{({\rm max\,Capa})}$, compared to FDSS without SE, as a function of FDSS shaping. The channel is TDL-C (300 ns) with $N_{\rm sc}=96$. Both FDSS window types are considered, and plain DFT-s-OFDM corresponds to an FDSS with 0dB attenuation.
For low data rate operating at low SNR, FDSS attenuation has little impact on the maximum achievable rate, and the SE size also has minimal impact. For higher rates operating at higher SNR, the effect of FDSS shaping and SE size becomes more pronounced. At 5dB SNR, the spectral efficiency loss from FDSS Hann windowing with 14-14dB power ripple attenuation, compared to plain DFT-s-OFDM without FDSS attenuation, is of $26\%$; however by using FDSS-SE with $N_{\rm e}^{({\rm max\,Capa})}$ for the same window, this loss can be reduced to $19\%$. The Kaiser windows outperforms the Hann window for strong shaping, which is becoming non-negligeable, especially in the higher-SNR regime. At 5dB SNR, the spectral efficiency loss for the Kaiser window with similar 14-14dB power attenuation is of $19\%$, which can be further reduced to $13\%$ with optimized SE size $N_{\rm e}^{({\rm max\,Capa})}$.
V-C4 Rate-PAPR Trade-Off
We show that they are two distinct optima for SE size: one for PAPR minimization ($N_{\rm e}^{({\rm min\,PAPR})}$), and another for rate maximization ($N_{\rm e}^{({\rm max\,Capa})}$), which implies the existence of a range of SE sizes that can improve both PAPR and rate simultaneously. Fig. 15 compares $N_{\rm e}^{({\rm min\,PAPR})}$ and $N_{\rm e}^{({\rm max\,Capa})}$ as functions of shaping for both window types with $N_{\rm sc}=96$. For $N_{\rm e}^{({\rm max\,Capa})}$, the TDL-C (300ns) channel is considered at SNR 5-5, 0, and 5 dB. The $N_{\rm e}^{({\rm min\,PAPR})}$ curves in Figs. 15(a) and  15(b) are obtained by minimizing the bound (13) for QPSK.
For both window types, the SE size optimal for PAPR minimization is larger than that for rate maximization, and the gap widens with SNR. In general, the Kaiser window requires smaller SE sizes for both criteria. Interestingly, SE improves the rate of FDSS (i.e. $N_{\rm e}^{({\rm max\,Capa})}>0$) only when the FDSS attenuation is sufficiently large.
Fig. 14 also reports the achievable rate when using $N_{\rm e}^{({\rm min\,PAPR})}$ as in Figs. 15(a) and  15(b). With little shaping, this is detrimental (a loss of $\approx 5$% with no FDSS compared to plain DFT-s-OFDM), but with stronger shaping, $N_{\rm e}^{({\rm min\,PAPR})}$ can increase the rate. At 14-14dB attenuation, the rate loss with Hann window is 21% compared to plain DFT-s-OFDM (vs. 26% with no SE, and 19% with optimal SE), while for the Kaiser window it is of 17% (vs. 19% with no SE, and 13% with optimal SE).
Similarly, while not optimal for PAPR, the SE size $N_{\rm e}^{({\rm max\,Capa})}$ can still provide significant PAPR reduction, as shown on Figs. 15(c) and 15(d), which give the $10^{-3}$ CCDF PAPR values using the SE size on Figs. 15(a) and  15(b). The PAPR reduction with $N_{\rm e}^{({\rm max\,Capa})}$ is nearly optimal for the Hann window at large attenuation, but weaker for the Kaiser window.
Overall, $N_{\rm e}^{({\rm max\,Capa})}$ and $N_{\rm e}^{({\rm min\,PAPR})}$ define, respectively, a lower and upper bound on the SE size to be considered for a practical PAPR–rate trade-off under a given configuration and scenario.
VI Conclusion
This paper presented a systematic performance evaluation of FDSS with spectrum extension (FDSS-SE) by employing single-parameter FDSS window families, encompassing various SE methods in the literature. For $\pi/2$-BPSK, the symmetrical double-side SE is optimal, whereas for higher-order QAM, the single-side SE performs better while optimality is achieved by another asymmetric SE. More importantly, we showed that there exist two distinct optima for the SE size: one for PAPR minimization and another for rate maximization, demonstrating the possibility of simultaneously improving both metrics. The optimum SE for PAPR performance depends mainly of the FDSS window and is largely invariant to bandwidth and constellation order for regular QAM. The optimum SE for rate maximization depends on the FDSS window and the SNR, while remaining also nearly invariant to bandwidth. This suggests that if SE is introduced on top of FDSS in 6G, it should be adaptively configured based on the user’s FDSS implementation, if proprietary, and the transmission link quality.
-A 3-tap Filters as Deformed Hann Window
Consider a 3-tap filter defined as f=[b, 1,b]f=[-b,\ 1,\ -b]. The corresponding un-normalized FDSS window is
$W_{\rm 3-tap}[k^{\prime}]$ $=$ $\sum_{m=0}^{2}f[m]e^{-{\rm j}\frac{2\pi}{N_{\rm sc}}k^{\prime}m}$ (34)
$=$ $e^{-{\rm j}\frac{2\pi}{N_{\rm sc}}k^{\prime}}\left(1-2b\cos\frac{2\pi}{N_{\rm sc}}k^{\prime}\right).$ (35)
The linear phase ramp $e^{-{\rm j}\frac{2\pi}{N_{\rm sc}}k^{\prime}}$ has only the effect of a circularly shift on the OFDM symbol so it has no impact on the PAPR and can be ignored. Then, using the change of variable, $2b=\frac{1-\beta}{1+\beta}$, we have a parametrization that quantifies directly the maximum ripple of the window as then $\frac{\min_{k}W_{\beta}[k^{\prime}]}{\max_{k}W_{\beta}[k^{\prime}]}=\beta$.
We get (9) by applying an half index shift $k=k^{\prime}+1/2$ for symmetry such that $W_{\beta}[k]=W_{\beta}[N_{\rm sc}-1-k]$; and a normalization following from from the geometric sum formula showing that for a constant $\alpha$ we have
$\sum_{k=0}^{N_{\rm sc}-1}\left(1-\alpha\cos\frac{2\pi}{N_{\rm sc}}k\right)^{2}=N_{\rm sc}\left(1+\frac{\alpha^{2}}{2}\right).$ (36)
-B PAPR Bounds
Noticing that that the average power of an OFDM symbol is $\mathbb{E}{|s[n]|^{2}}=\frac{N_{\rm sc}}{N_{\rm fft}}$, the PAPR simplifies to
${\rm PAPR}=\frac{N_{\rm fft}}{N_{\rm sc}}\max_{0\leq n<N_{\rm fft}}|s[n]|^{2}.$ (37)
Upper bound on $\max_{0\leq n<N_{\rm fft}}|s[n]|^{2}$ then follows by adapting [4, Th. 1] and [5, Th. 1] provided for $\phi$-BPSK and QPSK, respectively, to any constellation and shifts. The case considered therein corresponds to $L=0$, with a shift $L$ then $\angle\frac{p_{i}[n]}{p_{j}[n]}=-(i-j)2\pi\frac{L}{N_{\rm data}}+\angle\frac{p_{0}\left[n-\frac{N_{\rm fft}}{N_{\rm data}}i\right]}{p_{0}\left[n-\frac{N_{\rm fft}}{N_{\rm data}}j\right]}$ where the last term is evaluated in [4, Lem. 1].
The second general upper bound follows directly from the fact that $u_{i,j}^{\mathcal{C}}\leq 1$.
-C Proof of (17)
From (6), the phase difference between two neighboring pulses with indices $m$ and $(m+1)$ is
$\angle\frac{p_{m+1}[n]}{p_{m}[n]}=-\frac{2\pi L}{N_{\rm data}}+\angle\frac{p_{0}\left[n-\frac{N_{\rm fft}}{N_{\rm data}}(m+1)\right]}{p_{0}\left[n-\frac{N_{\rm fft}}{N_{\rm data}}m\right]}.$ (38)
Then, as shown in [4, Lem. 1] if the FDSS window is real and symmetric we have
$\angle\frac{p_{0}\left[n-\frac{N_{\rm fft}}{N_{\rm data}}(m+1)\right]}{p_{0}\left[n-\frac{N_{\rm fft}}{N_{\rm data}}m\right]}=-\frac{(N_{\rm sc}-1)}{N_{\rm data}}\pi+\{0\text{ or }\pi\}$ (39)
which can also be directly verified to hold true for the Dirichlet pulse (8). Therefore we get the pulsed phase difference
$\angle\frac{p_{m+1}[n]}{p_{m}[n]}$ $=$ $-\frac{2\pi}{N_{\rm data}}\left(L+\frac{(N_{\rm sc}-1)}{2}\right)\;({\rm mod}\,\pi)$ (40)
$=$ $-\frac{2\pi}{N_{\rm data}}\left(L+\frac{(N_{\rm e}-1)}{2}\right)\;({\rm mod}\,\pi).$
-D Proofs of Prop. 2
Expanding (31) and using a change of variable, we have
$r[m]$ $=$ $\frac{1}{\sqrt{N_{\rm data}}}\sum_{k=0}^{N_{\rm data}-1}\tilde{R}_{\rm eq}[k-L]e^{{\rm j}\frac{2\pi}{N_{\rm data}}km}$ (41)
$=$ $\frac{e^{{\rm j}\frac{2\pi}{N_{\rm data}}mL}}{\sqrt{N_{\rm data}}}\sum_{k=0}^{N_{\rm data}-1}\tilde{R}_{\rm eq}[k]e^{{\rm j}\frac{2\pi}{N_{\rm data}}km}.$ (42)
Expanding (29) we get
$\tilde{R}_{\rm eq}[k]=\frac{G[k]}{G[k]+1}X[k+L]+\frac{\tilde{Z}[k]}{G[k]+1}.$ (43)
Plugging (3) in (43), and (43) in (41), we can rewrite  (41) as
$r[m]$ $=$ $\sum_{n=0}^{N_{\rm data}-1}g_{m-n}x[n]+n[m]$ (44)
$=$ $g_{0}x[m]+{\rm ICI}[m]+n[m]$
where $g_{0}$ is the desired channel component, independent of $m$, ${\rm ICI}[m]=\sum_{\begin{subarray}{c}n=0\\ n\neq m\end{subarray}}^{N_{\rm data}-1}g_{m-n}x[n]$ with
$g_{a}=\frac{e^{{\rm j}\frac{2\pi}{N_{\rm data}}La}}{N_{\rm data}}\sum_{k=0}^{N_{\rm data}-1}\frac{G[k]}{G[k]+1}e^{{\rm j}2\pi\frac{ka}{N_{\rm data}}},$ (45)
and the processed noise is
$n[m]=\frac{e^{{\rm j}\frac{2\pi}{N_{\rm data}}Lm}}{\sqrt{N_{\rm data}}}\sum_{k=0}^{N_{\rm data}-1}\frac{\tilde{Z}[k]}{G[k]+1}e^{{\rm j}2\pi\frac{km}{N_{\rm data}}}.$ (46)
The desired signal power is directly given by $g_{0}^{2}$. Noticing that the equalized noise power is $\sigma_{\tilde{Z}}^{2}=G[k]\sigma_{Z}^{2}$, the power of the processed noise (46) becomes $\sigma_{n}^{2}=\frac{1}{N_{\rm data}}\sum_{k=0}^{N_{\rm data}-1}\frac{G[k]}{(G[k]+1)^{2}}$.
The noise-plus-interference power term can be computed as similarly done in [33] and simplified to $(g_{0}-g_{0}^{2})$. To verify this, the ICI power can be first simplified as
$\sigma_{\rm ICI}^{2}$ $=$ $\sum_{\begin{subarray}{c}n=0\\ n\neq m\end{subarray}}^{N_{\rm data}-1}|g_{m-n}|^{2}=\left(\sum_{n=0}^{N_{\rm data}-1}|g_{m-n}|^{2}\right)-g_{0}^{2}$ (47)
$=$ $\frac{1}{N_{\rm data}}\sum_{k=0}^{N_{\rm data}-1}\frac{G[k]^{2}}{(G[k]+1)^{2}}-g_{0}^{2}$
where the last equality follows by direct expansion of (45) and simplification by the exponential sum formula which gives $\sum_{n=0}^{N_{\rm data}-1}e^{{\rm j}2\pi\frac{na}{N_{\rm data}}}=0$ for any non-zero integer $a$. Then by direct combination we have
$\sigma_{\rm ICI}^{2}+\sigma_{n}^{2}$ $=$ $\frac{1}{N_{\rm data}}\sum_{k=0}^{N_{\rm data}-1}\frac{G[k]^{2}+G[k]}{(G[k]+1)^{2}}-g_{0}^{2}$ (48)
$=$ $\frac{1}{N_{\rm data}}\sum_{k=0}^{N_{\rm data}-1}\frac{G[k]}{G[k]+1}-g_{0}^{2}$
$=$ $g_{0}-g_{0}^{2}.$
The corresponding SINR (32) follows accordingly.
Using this SINR expression, which is independent of $m$, the achievable rate (33) is obtained as the sum-capacity of the DFT-s-OFDM subcarriers. This follows by treating the subcarriers as parallel channels, and bounding the mutual information of each to the one of a Gaussian additive noise channel with known noise variance, i.e. conditioned also on the knowledge of the effective interference channel gains. The rate is thus an achievable rate under the assumption of perfect knowledge of channel state information at the receiver.